OpenStudy (anonymous):

A square has an area of 16 cm². Calculate the length of its diagonal to the nearest 10th of a centimeter. A. 6.4 cm B. 5.7 cm C. 8 cm D. 4 centimeters E. 64 cm

OpenStudy (kainui):

Since it's a square you know that its length and width are the same, so the square root of your area is your length of each side. But to get your diagonal, you can see that it's just a right triangle, so use the pythagorean theorem to solve for the hypotenuse!

jhonyy9 (jhonyy9):

diagonal=sqrt32 =4sqrt2 = 5,7 cm so B. is right

jhonyy9 (jhonyy9):

so if area =16 =side *side = side ^2 = ---- sqrt16=length of side = 4 - diagonal ^2 = side ^2 +side ^2 - diagonal ^2 =16+16 =32 - diagonal = sqrt32 =sqrt(16*2)=4sqrt2=4*1,41=5,61 aprox. 5,7